a)How many grams are in 2.250 moles of N2?

b)How many grams are in 2.250 moles of NaBr?

c)How many grams are in 2.250 moles of C6H14?
Answer #1

Number of moles =mass / molar mass

Mass = (number of moles) ( molar mass)

A) mass of N2 = moles of N2 (molar mass of N2)

Molar mass of N2 = 28 g / mole

Mass of N2= 2.25 (28)= 63 grams

B) molar mass of NaBr 103 g/ mole

Mass of NaBr= 2.25 (103) =231.75 grams

C) molar mass of c6 H 14 =86 gram / mole

Mass of C6 H14 = 2.25 (86)= 193.5 grams